Status Report

NASA Space Station On-Orbit Status 25 June 2005

By SpaceRef Editor
June 26, 2005
Filed under , , ,
NASA Space Station On-Orbit Status 25 June 2005

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below. Saturday — off-duty day for Sergei Krikalev and John Phillips, except for some housekeeping and voluntary work.

The crew completed the regular weekly 3-hr task of thorough station cleaning, wearing protective garment. [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other surfaces and the CDR’s sleep station with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

FE/SO Phillips collected the periodic reading of the cabin air’s current CO2 partial pressure in the SM and Lab, using the U.S. CDMK (carbon dioxide monitoring kit, #1013), for calldown, along with the battery status, for use in trending analyses.

The CDR did the daily routine maintenance of the Service Module (SM)’s SOZh environment control & life support system, including its ASU toilet system and today also the weekly inspection of the BRPK air/liquid condensate separator apparatus.

John performed the regular weekly maintenance reboot on the operational PCS (Portable Computer System) laptops.

Processing Status
Daily Mission
Return to Flight
Weekly Status
Weekly Science
Daily On-Orbit Status
Daily Crew Timeline
Soyuz | Progress
ISS News | ATV

On Sergei’s task list for today was his third run of the regular monthly sessions of the ETD experiment (=Eye Tracking Device, to investigate the coordination of eye and head movements in zero-G), taking place in the DC-1 docking module’s center sphere. [After a calibration with the calibrating unit, the experiment investigates horizontal eye and head movement coordination, measured Listing’s plane, and determined the orientation of the vestibulo-ocular coordinate system, using five target marks on the horizontal plane. Each step required another prior calibration run, using visual target cues or the calibration unit.]

Also added to the discretionary “job jar” task list for Krikalev was another periodic temperature check with the automatic temperature recorder (ART) on the BIO-11″Statokonia” payload and its ULITKA (“snail”) incubator, set up in the SM since 3/3 and replaced with new gear delivered on 18P. [BIO-11 studies the composition of statoconia, i.e., the organ of equilibrium in snails, and other phenomena exhibited by “ulitka” in zero-G and post-flight.]

A third “job jar” task for the CDR is to complete gathering & readying of all equipment items required for next week’s (6/27-28) installation and set-up of the ATV (Automated Transfer Vehicle) control panel (BUAP), antenna feeder unit and space-to-space radio (MBRL). Extensive testing of this ATV PCE (proximity communications equipment) will then follow on 6/28-30.

At ~6:30am, the crew held their weekly teleconference with ISS Program Management at JSC/Houston via S-band/audio.

At ~9:40am EDT, the crew conducted the regular weekly planning conference (WPC) with the ground, discussing next week’s “Look-Ahead Plan” (prepared jointly by MCC-H and TsUP/Moscow timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.

Both crewmembers worked out in their regular 2.5-hr. physical exercise program on the TVIS treadmill, RED resistive exerciser and VELO ergometer bike with bungee cord load trainer. [Sergei’s daily protocol prescribes a strict four-day microcycle exercise with 1.5 hr on the treadmill and one hour on VELO plus load trainer (today: Day 1 of a new set).]

Afterwards, the FE transferred the exercise data files to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

The crew replaced the SM PCS (Portable Computer System) battery, which wasn’t holding its charge, with a fully charged battery from the ER3 (Express Rack #3) laptop which currently is not in use. A plan is in work to identify all available laptop batteries and their current SOC (state of charge).

Working off his “job jar” task list, Sergei conducted the regular daily inspection of the BIO-5 Rasteniya-2 (“Plants-2”) experiment, including filling its water canister for the Lada-7 greenhouse as required.

On John’s task list, MCC-H has added the task to locate a couple of good candidate batteries for 760XD laptops. [One of these batteries is to be installed in the PCS (portable computer system) shell currently in the Airlock. The other battery should be stowed in a convenient location for future use.]

Also on the “job jar” task list is the troubleshooting of the RPC-17 (remote power controller #17) of the Node UOP-2 (utility outlet panel 2) that tripped yesterday during John’s testing of the CBCS (centerline berthing camera system) for the MPLM Raffaello docking. [Data dumped to the ground showed that the cause of the trip was an over-current, which is similar to what the Expedition 10 crew saw back in November on the Lab UOP-4. The crew will therefore go through the same troubleshooting procedure used by Chiao and Sharipov.]

Weekly Science Update (Expedition Eleven — 9th):

Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology (HRF GASMAP): Next week the crew will perform the GASMAP Functional Check, which is done once per Increment.

Human Research Facility/Workstation (HRF WS): Continuing.

Advanced Ultrasound (ADUM): Continuing.

Renal Stone (RS): Next data collection in July.

Foot/Ground Reaction Forces during Space Flight (FOOT): The data from last week’s session indicate that the Foot data collection run was successful. Corrective comments for data collection #3 have been uplinked to Dr. Phillips, to ensure another successful data run…

Space Acceleration Measurement System (SAMS): SAMS is nominal and receiving acceleration data.

Microgravity Acceleration Measurement System (MAMS): MAMS remains in nominal operations.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES): PCG-STES is performing nominally.

Binary Colloidal Alloy Test-3 (BCAT-3): BCAT-3 Slow Growth Sample Module will be left undisturbed in its current location by the E11 crew. In order for the samples to potentially grow crystals that can be photographed during Increment 12 operations, the Sample Module must be left undisturbed.

Materials ISS Experiment (MISSE): In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data. To be exchanged during LF-1.

Dust and Aerosol Measurement Feasibility Test (DAFT): Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM): The team is looking forward to the next session in July.

Serial Network Flow Monitor (SNFM): Completed.

Fluid Merging Viscosity Measurement (FMVM): Nothing new.

Space Experiment Module (SEM): Nothing new. Experimenters and kids are working to get the next two satchels on ULF1.1.

Miscible Fluids in Microgravity (MFMG): MFMG payload operations are finished.

Educational Payload Operations (EPO): in progress.

Crew Earth Observations (CEO): As of 6/20 the ground has received and reviewed a total of 3,424 of ISS CEO images. An excellent detailed image of Sept-Iles, Quebec is featured on NASA/GSFC’s Earth Observatory website this weekend. To crew: “Camera times look good now. Please continue your practice with the long lenses in nadir views of familiar, high-contrast features.”

Today’s CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Glacial features, Western Libya (meandering subglacial river beds have been re-exposed by erosion since they were first formed ~400 million years ago. ISS/CEO imagery should allow the regional pattern to be detected. Nadir views were requested for ~30 secs), Angola fires (Dynamic event. Looking mainly left into the dense savanna woodlands that lie south of the Congo Rainforest. These savannas burn more often than any other woodland on Earth), Andean dust, Argentina (Dynamic event. Mid-winter is high season for dust plumes blowing east off the high central Andes. There is a hint these may be operating now. The special interest of the dust storms is that they show an unexpected source for soil nutrient for the Amazon and other lowlands east of the Andes. Looking left and right), and Wake Island reefs, NW Pacific (nadir pass. Detailed images for mapping of the coral reef were requested).

CEO photography can be viewed and studied at the websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 11 crew visit:

Expedition 11 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Some Increment 11 Main Events (not final):

  • Reboost — 6/29 (4:21pm, delta-V 2.3 m/s);
  • LF-1/STS-114 launch — 7/13 (18-day window opens);
  • LF-1/STS-114 dock — 7/15 (adding 110,713 kg to ISS mass);
  • LF-1/STS-114 undock — 7/23;
  • Soyuz TMA-6/10S relocate (from DC-1 to FGB) — ~8/16;
  • Progress M-54/19P launch – TBD;
  • Progress M-53/18P undock — TBD;
  • ULF1.1/STS-121 launch — NET 9/9 (launch window opens);
  • ULF1.1/STS-121 dock — 9/11;
  • ULF1.1/STS-121 undock — 9/19;
  • Soyuz TMA-7/11S launch — 9/27;
  • Soyuz TMA-7/11S dock — 9/29;
  • 12A/STS-115 launch — NET 2/16/06;
  • Soyuz TMA-7/11S relocate (from DC-1 to FGB) — ~10/15;
  • 12A.1/STS-116 launch — NET 4/23/06;
  • 13A/STS-117 launch — NET 7/13/06.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit (as of this morning, 8:20am EDT [= epoch]):

  • Mean altitude — 350.0 km
  • Apogee height — 352.7 km
  • Perigee height — 347.3 km
  • Period — 91.54 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004053
  • Solar Beta Angle — 55.8 deg (magnitude increasing)
  • Orbits per 24-hr. day — 15.73
  • Mean altitude loss in last 24 hours — 74
  • Revolutions since FGB/Zarya launch (Nov. 98) — 37699

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.